A manufacturer of chocolate chips would like to know whether its bag filling machine works correctly at the 430 gram setting. It is believed that the machine is overfilling the bags. A 37 bag sample had a mean of 439 grams. Assume a population standard deviation of 29. Is there sufficient evidence at the 0.05 level that the bags are overfilled? 24. Step 1. State the hypotheses: Ho: Step 2. Find the value of the z test statistic. (Round your answer to 2 decimal places.)
media%2F943%2F943252df-6723-4f5a-bdb6-b6
Answer #1

The hypotheses are

H_0:\mu \leqslant 430\\ H_1:\mu >430\\

The \left ( 1-\alpha \right )100\% CI for mean is \overline{x}\pm z_{1-\alpha }\frac{\sigma }{\sqrt{n}} .

The test statistic is

z=\frac{\overline{x}-430}{\frac{\sigma }{\sqrt{n}}}\\ z=\frac{439-430}{29/\sqrt{37}}\\ {\color{Blue} z=1.888}

This is a right tailed test. Hence One-tailed test

The P-value is

\textup{P-value}=P\left ( z >1.888\right )\\ \textup{P-value}=1-\Phi \left ( 1.888\right )\\ {\color{Blue} \textup{P-value}= 0.0295}

The level of significance is \alpha =0.05 .

Since \textup{P-value}= 0.0295<0.05=\alpha ,

We Reject the null Hypothesis. (A).